OpenStudy (anonymous):

derivative of (1+cos^2 7x)^3

OpenStudy (anonymous):

help please!

OpenStudy (asnaseer):

\[f(x) = (1 + \cos ^{2}7x) ^{3}\]let \[y = 1 + \cos ^{2}7x\]so \[dy/dx = (2\cos7x)(-\sin7x)(7) = -14\cos7x\sin7x\] \[f(x) = y ^{3}\] \[f\prime(x) = 3y ^{2}dy/dx = 3y ^{2}(-14\cos7x\sin7x) = -42y ^{2}\cos7x\sin7x\] \[f\prime(x) = -42(1+\cos ^{2}7x) ^{2}\cos7x\sin7x\]